OpenStudy (anonymous):

What is the fifth term of the geometric sequence? a1 = 120, a2 = 36, a3 = 10.8, a6 = 0.2916

OpenStudy (smored):

By looking at the sequence, i can guess that it is a division (difference is not constant). 120/36=10/3 \[\frac{ 120 }{ \frac{ 10 }{ 3 } }=36\] \[\frac{ 36 }{ \frac{ 10 }{ 3 } }=10.8\] \[\frac{ 10.8 }{ \frac{ 10 }{ 3 } }=0.2916\] \[\frac{ 0.2916 }{ \frac{ 10 }{ 3 } }=0.08748\] so the answer is 0.08748

OpenStudy (anonymous):

a_n = a_1 * r^{n-1} 36/120 = 120/120 3/10 = r a_6=a_n*r 10/3 .2916 = 1)5 * 3/10 * 10/3 .972 = a_5

OpenStudy (ranga):

36/120 = 0.3 The common ratio is 0.3 a5 * 0.3 = a6 a5 = a6/0.3 = 0.2916/0.3 = 0.972